find an equaction of a line with x-intercept 6 and y-intercept -5.
Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

find an equaction of a line with x-intercept 6 and y-intercept -5.
-----------
With x-int of A and y-int of B:
Bx + Ay = A*B

Answer by rothauserc(4718)

using slope-intercept form, we have
:
y = mx -5, where m is the slope of the line
:
this tells us that when x is 0, the y intercept is -5
:
now solve this equation for x
:
x = (y+5) / m
:
we know that the x intercept is 6, therefore
:
6 = (0+5) / m
:
m = 5/6
:
*************************************************
y = (5/6)x -5 (slope intercept form of the line)
*************************************************
:
the standard form of a line is
:
Ax +By = C, where A is a positive integer and B and C are integers
:
Beginning with the slope intercept form, subtract y from both sides of =
:
(5/6)x -y -5 = 0
:
(5/6)x -y = 5
:
now multiply both sides of = by 6
:
**************************************************
5x -6y = 30 (standard form of the line)
**************************************************